# Generalizable, Fast, and Accurate DeepQSPR with fastprop

### v4

@@ Abstract
Quantitative Structure-Property Relationship studies (QSPR), often referred to interchangeably as QSAR, seek to establish a mapping between molecular structure and an arbitrary Quantity of Interest (QOI). Historically this was done on a QOI-by-QOI basis with new descriptors being devised by researchers to specifically map to their QOI. A large number of descriptors have been invented, and can be computed using packages like DRAGON (later E-dragon), PaDEL-descriptor (and padelpy), Mordred, CODESSA, and many others. The sheer number of different descriptor packages resulted in the creation of ‘meta-packages’ which served only to aggregate these other calculators, including tools like molfeat, ChemDes, Parameter Client, and AIMSim.
Generalizable descriptor-based modeling was a natural evolution of these meta-packages’ development. Historically QSPR researchers focused almost exclusively on linear methods. Another community of researchers focused on finding nonlinear correlations between molecular structures and a QOI, often using Deep learning (DL). The DL community typically used molecular fingerprints instead of the complex descriptors popular in QSPR community. Recently the DL community has turned to learned representations primarily via message passing graph neural networks. This approach has proved remarkably effective but is not without drawbacks. Learning a representation requires large datasets to avoid over-fitting or even learn at all, loses interpretability since an embedding’s meaning can only be induced retroactively, and needs significant execution time given the complexity of the underlying message passing algorithm. This paper introduces fastprop, a software package and general Deep-QSPR framework that combines a cogent set of molecular descriptors with DL to achieve state-of-the-art performance on datasets ranging from tens to tens of thousands of molecules. fastprop is designed with Research Software Engineering best practices and is free and open source, hosted at github.com/jacksonburns/fastprop.
@@ Introduction
Early in the history of computing, limited computational power meant that significant human expertise was required to guide QSPR models toward effectiveness. This materialized in the form of bespoke molecular descriptors: the Wiener Index in 1947 [2], Atom-Bond Connectivity indices in 1998 [3], and thousands of others. To this day descriptors are still being developed - the geometric-harmonic-Zagreb degree based descriptors were proposed by Arockiaraj et al. in 2023 [4]. In each case, domain experts devised an algorithm which mapped a molecular structure to some scalar value. This algorithm would take into account features of the molecule which that expert deduced were relevant to the property at hand. This time consuming technique is of course highly effective but the dispersed nature of this chemical knowledge means that these descriptors are spread out throughout many journals and domains with no single source to compute them all.
@@
Efforts are of course underway to address this limitation, though none are broadly successful. One simple but incredibly computationally expensive approach is to use delta learning, which artificially increases dataset size by generating all possible pairs of molecules from the available data (thus squaring the size of the dataset). This was attempted by Nalini et al. [15], who used an unmodified version of Chemprop referred to as ‘DeepDelta’ to predict differences in molecular properties for pairs of molecules. They achieve increased performance over standard LR approaches but lost the ability to train on large datasets due to simple runtime limitations. Other increasingly complex approaches are discussed in the outstanding review by van Tilborg et al. [16].
@@
At its core the fastprop ‘architecture’ is simply the mordred molecular descriptor calculator 222The original mordred package is no longer maintained. fastprop uses a fork of mordred called mordredcommunity that is maintained by community-contributed patches (see github.com/JacksonBurns/mordred-community). Multiple descriptor calculators from the very thorough review by McGibbon et al. [18] could be used instead, though none are as readily interoperable as mordred. [17] connected to a Feedforward Neural Network (FNN) implemented in PyTorch Lightning [19] (Figure 1) - an existing approach formalized into an easy-to-use, reliable, and correct implementation. fastprop is highly modular for seamless integration into existing workflows and includes end-to-end interfaces for general use. In the latter mode the user simply specifies a set of SMILES [20], a linear textual encoding of molecules, and their corresponding properties. fastprop automatically calculates and caches the corresponding molecular descriptors with mordred, re-scales both the descriptors and the targets appropriately, and then trains an FNN to predict the indicated target properties. By default this FNN is two hidden layers with 1800 neurons each connected by ReLU activation functions, though the configuration can be readily changed via the command line interface or configuration file. fastprop owes its success to the cogent set of descriptors assembled by the developers of mordred, the ease of training FNNs with modern software like PyTorch Lightning, and the careful application of Research Software Engineering best practices that make it as user friendly as the best-maintained alternatives.
@@
The reason fastprop stands out from these studies and contradicts previous reports is for the simple reason that it works. As discussed at length in the Results & Discussion section, this approach matches the performance of leading LR approaches on common benchmark datasets and bespoke QSPR models on small real-world datasets. fastprop also overcomes the limitations of LRs discussed above. Because all inputs to the FNN are physically meaningful molecular descriptors, intermediate representations in the FNN are also physically meaningful and can be directly interpreted. The simplicity of the framework enables domain experts to apply it easily and makes model training dramatically faster than LRs. Most importantly this approach is successful on the smallest of real-world datasets. By starting from such an informed initialization the FNN can be readily trained on datasets with as few as forty training examples (see PAHs).
@@
The method for splitting data into training, validation, and testing sets varies on a per-study basis and is described in each sub-section. Sampling is performed using the astartes package [28] which implements a variety of sampling algorithms and is highly reproducible. For datasets containing missing target values or invalid SMILES strings, those entries were dropped.
@@
and classification: - Area Under the Receiver Operating Curve (AUROC, AUC, or ROC-AUC): Summary statistic combining all possible classification errors; scale-independent, range 0.5 (worst, random guessing) to 1.0 (perfect classifier). - Accuracy: Fraction of correct classifications, expressed as either a percentage or a number; scale-independent, range 0 (worst) to 1 (perfect classifier).
@@
See Table LABEL:results_table for a summary of all the results. Subsequent sections explore each in greater detail.
Benchmark | Samples (k) | Metric |
@@
QM9 | ~134 | MAE | 0.0047a | 0.0060 | 0.0081a | ~ |
Pgp | ~1.3 | AUROC | 0.94b | 0.90 | 0.89b | ~ |
ARA | ~0.8 | Accuracy | 91c | 88 | 82* | 0.083 |
Flash | ~0.6 | RMSE | 13.2d | 13.0 | 21.2* | 0.021 |
YSI | ~0.4 | MAE | 22.3e | 25.0 | 28.9* | 0.29 |
PAH | ~0.06 | R2 | 0.96f | 0.97 | 0.59* | 0.0012 |
@@
Statistical comparisons of fastprop to Chemprop (shown in the p column) are performed using the non-parametric Wilcoxon-Mann-Whitney Test as implemented in GNumeric. Values are only shown for results generated in this study which are known to be performed using the same methods. Only the results for Flash and PAH are statistically significant at 95% confidence (p<0.05).
@@
Originally described in Scientific Data [27] and perhaps the most established property prediction benchmark, Quantum Machine 9 (QM9) provides quantum mechanics derived descriptors for many small molecules containing one to nine heavy atoms, totaling ~134k. The data was retrieved from MoleculeNet [25] in a readily usable format. As a point of comparison, performance metrics are retrieved from the paper presenting the UniMol architecture [10] previously mentioned. In that study they trained on only three especially difficult QOIs (homo, lumo, and gap) using scaffold-based splitting (a more challenging alternative to random splitting), reporting mean and standard deviation across 3 repetitions.
@@
First reported in 2011 by Broccatelli and coworkers [33], this dataset has since become a standard benchmark and is included in the Therapeutic Data Commons (TDC) [34] model benchmarking suite. The dataset maps approximately 1.2k small molecule drugs to a binary label indicating if they inhibit P-glycoprotein (Pgp). TDC serves this data through a Python package, but due to installation issues the data was retrieved from the original study instead. The recommended splitting approach is a 70/10/20 scaffold-based split which is done here with 4 replicates.
@@
Compiled by Schaduangrat et al. in 2023 [30], this dataset maps ~0.8k small molecules to a binary label indicating if the molecule is an Androgen Receptor Antagonist (ARA). The reference study introduced DeepAR, a highly complex modeling approach, which achieved an accuracy of 0.911 and an AUROC of 0.945.
For this study an 80/10/10 random splitting is repeated four times on the dataset since no analogous split to the reference study can be determined. Chemprop takes 16 minutes and 55 seconds to run on this dataset and achieves only 0.824 0.020 accuracy and 0.898 0.022 AUROC. fastprop takes only 1 minute and 54 seconds (1 minute and 39 seconds for descriptor calculation) and is competitive with the reference study in performance, achieving a 88.2 3.7% accuracy and 0.935 0.034 AUROC. The purely linear QSPR model falls far behind these methods with a 71.8 6.6% accuracy and 0.824 0.052 AUROC.
@@
First assembled and fitted to by Saldana and coauthors [31] the dataset (Flash) includes around 0.6k entries, primarily alkanes and some oxygen-containing compounds, and their literature-reported flash point. The reference study reports the performance on only one repetition, but manually confirms that the distribution of points in the three splits follows the parent dataset. The split itself was a 70/20/10 random split, which is repeated four times for this study.
@@
Assembled by Das and coauthors [32] from a collection of other smaller datasets, this dataset maps ~0.4k molecular structures to a unified-scale Yield Sooting Index (YSI), a molecular property of interest to the combustion community. The reference study performs leave-one-out cross validation to fit a per-fragment contribution model, effectively a training size of >99%, without a holdout set. Though this is not standard practice and can lead to overly optimistic reported performance, the results will be carried forward regardless. The original study did not report overall performance metrics, so they have been re-calculated for this study using the predictions made by the reference model as provided on GitHub 444Predictions are available at this permalink to the CSV file on GitHub.. For comparison fastprop and Chemprop use a more typical 60/20/20 random split and 8 repetitions. Results are summarized in Table LABEL:ysi_results_table.
@@
DeepDelta reported an RMSE of 0.830 0.023 at predicting differences, whereas a typical Chemprop model trained to directly predict property values was only able to reach an accuracy of 0.965 0.019 when evaluated on its capacity to predict property differences. fastprop is able to outperform Chemprop, though not DeepDelta, achieving an RMSE of 0.930 0.029 when using the same splitting procedure above. It is evident that delta learning is still a powerful technique for regressing small datasets.
For completeness, the performance of Chemprop and fastprop on Fubrain are also compared to the original study. The study that first generated this dataset used mordred descriptors but as is convention they strictly applied linear modeling methods. Using both cross validation and and external test sets, they had an effective training/validation/testing split of 0.64/0.07/0.28 which will be repeated 4 times here for comparison. All told, their model achieved an RMSE of 0.53 averaged across all testing data. In only 39 seconds, of which 31 are spent calculating descriptors, fastprop far exceeds the reference model with an RMSE of 0.207 0.024. This also surpasses Chemprop, itself outperforming the reference model with an RMSE of 0.223 0.036.
@@
Though not discussed here for the sake of length, fastprop already contains the functionality to perform feature importance studies on trained models. By using SHAP values [38] to assign a scalar ‘importance’ to each of the input features, users can determine which of the mordred descriptors has the largest impact on model predictions. The utility of these values can be explored in greater detail on a case-by-case basis.
@@ Acknowledgements
pre 4. Arockiaraj M, Paul D, Clement J, et al (2023) Novel molecular hybrid geometric-harmonic-zagreb degree based descriptors and their efficacy in QSPR studies of polycyclic aromatic hydrocarbons. SAR and QSAR in Environmental Research 34:569–589. https://doi.org/10.1080/1062936x.2023.2239149
@@
pre 5. Ma J, Sheridan RP, Liaw A, et al (2015) Deep neural nets as a method for quantitative structure-activity relationships. Journal of Chemical Information and Modeling 55:263–274. https://doi.org/10.1021/ci500747n
@@
pre 6. Coley CW, Barzilay R, Jaakkola TS, et al (2017) Prediction of organic reaction outcomes using machine learning. ACS Central Science 3:434–443. https://doi.org/10.1021/acscentsci.7b00064
@@
pre 7. Yang K, Swanson K, Jin W, et al (2019) Analyzing learned molecular representations for property prediction. Journal of Chemical Information and Modeling 59:3370–3388. https://doi.org/10.1021/acs.jcim.9b00237
@@
pre 8. Heid E, Greenman KP, Chung Y, et al (2024) Chemprop: A machine learning package for chemical property prediction. Journal of Chemical Information and Modeling 64:9–17. https://doi.org/10.1021/acs.jcim.3c01250
@@
pre 9. Song Y, Zheng S, Niu Z, et al (2021) Communicative representation learning on attributed molecular graphs. In: Proceedings of the twenty-ninth international joint conference on artificial intelligence. Yokohama, Yokohama, Japan
@@
pre 10. Zhou G, Gao Z, Ding Q, et al (2023) Uni-mol: A universal 3D molecular representation learning framework. In: The eleventh international conference on learning representations
@@
pre 11. Chen J, Schwaller P (2023) Molecular hypergraph neural networks. https://doi.org/10.48550/arXiv.2312.13136
@@
pre 12. Zhao B, Xu W, Guan J, Zhou S (2023) Molecular property prediction based on graph structure learning. arXiv preprint arXiv:231216855
@@
pre 13. Wang Z, Jiang T, Wang J, Xuan Q (2024) Multi-modal representation learning for molecular property prediction: Sequence, graph, geometry. https://doi.org/10.48550/ARXIV.2401.03369
@@
pre 14. Zhu Y, Chen D, Du Y, et al (2024) Molecular contrastive pretraining with collaborative featurizations. Journal of Chemical Information and Modeling 64:1112–1122. https://doi.org/10.1021/acs.jcim.3c01468
@@
pre 15. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
@@
pre 16. Tilborg D van, Brinkmann H, Criscuolo E, et al (2024) Deep learning for low-data drug discovery: Hurdles and opportunities. ChemRxiv. https://doi.org/10.26434/chemrxiv-2024-w0wvl
@@
pre 17. Moriwaki H, Tian Y-S, Kawashita N, Takagi T (2018) Mordred: A molecular descriptor calculator. Journal of Cheminformatics 10:4. https://doi.org/10.1186/s13321-018-0258-y
pre 18. McGibbon M, Shave S, Dong J, et al (2023) From intuition to AI: evolution of small molecule representations in drug discovery. Briefings in Bioinformatics 25:bbad422. https://doi.org/10.1093/bib/bbad422
pre 19. Falcon W, The PyTorch Lightning team (2019) PyTorch Lightning
@@
pre 20. Weininger D (1988) SMILES, a chemical language and information system. 1. Introduction to methodology and encoding rules. Journal of Chemical Information and Computer Sciences 28:31–36. https://doi.org/10.1021/ci00057a005
@@
pre 21. Comesana AE, Huntington TT, Scown CD, et al (2022) A systematic method for selecting molecular descriptors as features when training models for predicting physiochemical properties. Fuel 321:123836. https://doi.org/https://doi.org/10.1016/j.fuel.2022.123836
@@
pre 22. Wu J, Wang S, Zhou L, et al (2020) Deep-learning architecture in QSPR modeling for the prediction of energy conversion efficiency of solar cells. Industrial and Engineering Chemistry Research 59:18991–19000. https://doi.org/10.1021/acs.iecr.0c03880
@@
pre 23. Esaki T, Ohashi R, Watanabe R, et al (2019) Computational model to predict the fraction of unbound drug in the brain. Journal of Chemical Information and Modeling 59:3251–3261. https://doi.org/10.1021/acs.jcim.9b00180

### v5

@@ Abstract
Quantitative Structure-Property Relationship studies (QSPR), often referred to interchangeably as QSAR, seek to establish a mapping between molecular structure and an arbitrary target property. Historically this was done on a target-by-target basis with new descriptors being devised to specifically map to a given target. Today software packages exist that calculate thousands of these descriptors, enabling general modeling typically with classical and machine learning methods. Also present today are learned representation methods in which deep learning models generate a target-specific representation during training. The former requires less training data and offers improved speed and interpretability while the latter offers excellent generality, while the intersection of the two remains under-explored.
This paper introduces fastprop, a software package and general Deep-QSPR framework that combines a cogent set of molecular descriptors with deep learning to achieve state-of-the-art performance on datasets ranging from tens to tens of thousands of molecules. fastprop provides both a user-friendly Command Line Interface and highly interoperable set of Python modules for the training and deployment of feedforward neural networks for property prediction. This approach yields improvements in speed and interpretability over existing methods while statistically equaling or exceeding their performance across most of the tested benchmarks. fastprop is designed with Research Software Engineering best practices and is free and open source, hosted at github.com/jacksonburns/fastprop.
@@ Introduction
Early in the history of computing, limited computational power meant that significant human expertise was required to guide QSPR models toward effectiveness. This materialized in the form of bespoke molecular descriptors - scalar-valued functions which operate on the molecular graph in such a way to reflect relevant structural and electronic information. Examples include rudimentary counting descriptors, the Wiener Index in 1947 [2], Atom-Bond Connectivity indices in 1998 [3], and many others [4]. To this day descriptors are still being developed - the geometric-harmonic-Zagreb degree based descriptors were proposed by Arockiaraj et al. in 2023 [5]. This time consuming technique is of course highly effective but the dispersed nature of this chemical knowledge means that these descriptors are spread out throughout many journals and domains with no single source to compute them all.
@@
Efforts are of course underway to address this limitation, though none are broadly successful. One simple but incredibly computationally expensive approach is to use delta learning, which artificially increases dataset size by generating all possible pairs of molecules from the available data (thus squaring the size of the dataset). This was attempted by Nalini et al. [16], who used an unmodified version of Chemprop referred to as ‘DeepDelta’ to predict differences in molecular properties for pairs of molecules. They achieve increased performance over standard LR approaches but lost the ability to train on large datasets due to simple runtime limitations. Another promising line of inquiry is the Transformer-CNN model of Karpov et al. [17], which leverages a pre-trained transformer model for prediction, circumventing the need for massive datasets and offering additional benefits in interpretability. This model is unique in that it operates directly on the SMILES representation of the molecule, also offering benefits in structural attribution of predictions. Other increasingly complex approaches are discussed in the outstanding review by van Tilborg et al. [18].
@@
At its core the fastprop ‘architecture’ is simply the mordred molecular descriptor calculator 222The original mordred package is no longer maintained. fastprop uses a fork of mordred called mordredcommunity that is maintained by community-contributed patches (see github.com/JacksonBurns/mordred-community). [19] connected to a Feedforward Neural Network (FNN) implemented in PyTorch Lightning [20] (Figure 1) - an existing approach formalized into an easy-to-use, reliable, and correct implementation. fastprop is highly modular for seamless integration into existing workflows and includes and end-to-end Command Line Interface (CLI) for general use. In the latter mode the user simply specifies a set of SMILES [21], a linear textual encoding of molecules, and their corresponding target values. fastprop optionally standardizes input molecule and then automatically calculates and caches the corresponding molecular descriptors with mordred, re-scales both the descriptors and the targets appropriately, and then trains an FNN to predict the indicated targets. By default this FNN is two hidden layers with 1800 neurons each connected by ReLU activation functions, though the configuration can be readily changed via the CLI or configuration file. Multitask regression and multi-label classification are also supported and configurable in the same manner, the former having been shown to significantly improve predictive power in cheminformatics models [22]. fastprop principally owes its success to the cogent set of descriptors assembled by the developers of mordred. Multiple descriptor calculators from the very thorough review by McGibbon et al. [23] could be used instead, though none are as readily interoperable as mordred. Additionally, the ease of training FNNs with modern software like PyTorch Lightning and the careful application of Research Software Engineering best practices make fastprop as user friendly as the best-maintained alternatives.
@@
The reason fastprop stands out from these studies and contradicts previous reports is for the simple reason that it works. As discussed at length in the Results & Discussion section, this approach statistically matches or exceeds the performance of leading LR approaches on common benchmark datasets and bespoke QSPR models on small real-world datasets. fastprop also overcomes the limitations of LRs discussed above. The FNN architecture and use of physically meaningful molecular descriptors enables the application of SHAP [28], a common tool for feature importance analysis (see Interpretability). The simplicity of the framework enables domain experts to apply it easily and makes model training dramatically faster than LRs. Most importantly this approach is successful on the smallest of real-world datasets. By starting from such an informed initialization the FNN can be readily trained on datasets with as few as forty training examples (see PAHs).
@@
The method for splitting data into training, validation, and testing sets varies on a per-study basis and is described in each sub-section. Sampling is performed using the astartes package [32] which implements a variety of sampling algorithms and is highly reproducible. For datasets containing missing target values or invalid SMILES strings, those entries were dropped, as is the default behavior of fastprop.
@@
Because the diversity of methods across these different datasets complicates inter-dataset comparisons, an additional set of benchmarks using an identical method across all datasets is included in Additional File 1, Table S1. This benchmark compares fastprop, Chemprop, and the aforementioned Transformer-CNN, which is especially suitable for the small datasets included therein. No long-form commentary on benchmark results is provided, though the conclusions are largely the same as those shown here.
@@
and classification: - Area Under the Receiver Operating Curve (AUROC, AUC, or ROC-AUC): Summary statistic combining all possible classification errors; scale-independent, range 0.5 (worst, random guessing) to 1.0 (perfect classifier). - Accuracy: Fraction of correct classifications, expressed as a percentage; scale-independent, range 0 (worst) to 100 (perfect classifier).
@@
See Table LABEL:results_table for a summary of all the results. Subsequent sections explore each in greater detail. For benchmarks with statistics, practically significant best performers are shown in bold.
Benchmark | Samples | Metric |
@@
QM9 | 133,885 | MAE | 0.0047a | 0.0060 | 0.0081a | ~ |
Pgp | 1,275 | AUROC | 0.94b | 0.90 | 0.89b | ~ |
ARA | 842 | Accuracy | 91c | 88 | 82* | 0.083 |
Flash | 632 | RMSE | 13.2d | 13.0 | 21.2* | 0.021 |
YSI | 442 | MAE | 22.3e | 25.0 | 28.9* | 0.29 |
PAH | 55 | R2 | 0.96f | 0.97 | 0.59* | 0.0012 |
@@
Statistical comparisons of fastprop to Chemprop (shown in the p column) are performed using the non-parametric Wilcoxon-Mann-Whitney Test as implemented in GNumeric. Values are only shown for results generated in this study which are known to be performed using the same methods. Only the results for Flash and PAH are statistically significant at 95% confidence (p<0.05), see benchmark-specific subsections for confidence intervals.
@@
Originally described in Scientific Data [31] and perhaps the most established property prediction benchmark, Quantum Machine 9 (QM9) provides quantum mechanics derived descriptors for many small molecules containing one to nine heavy atoms, totaling 133,885. The data was retrieved from MoleculeNet [29] in a readily usable format. As a point of comparison, performance metrics are retrieved from the paper presenting the UniMol architecture [11] previously mentioned. In that study they trained on only three especially difficult targets (homo, lumo, and gap) using scaffold-based splitting (a more challenging alternative to random splitting), reporting mean and standard deviation across 3 repetitions.
@@
First reported in 2011 by Broccatelli and coworkers [37], this dataset has since become a standard benchmark and is included in the Therapeutic Data Commons (TDC) [38] model benchmarking suite. The dataset maps 1,275 small molecule drugs to a binary label indicating if they inhibit P-glycoprotein (Pgp). TDC serves this data through a Python package, but due to installation issues the data was retrieved from the original study instead. The recommended splitting approach is a 70/10/20 scaffold-based split which is done here with 4 replicates.
@@
Compiled by Schaduangrat et al. in 2023 [34], this dataset maps 842 small molecules to a binary label indicating if the molecule is an Androgen Receptor Antagonist (ARA). The reference study introduced DeepAR, a highly complex modeling approach, which achieved an accuracy of 91.1% and an AUROC of 0.945.
For this study an 80/10/10 random splitting is repeated four times on the dataset since no analogous split to the reference study can be determined. Chemprop takes 16 minutes and 55 seconds to run on this dataset and achieves only 82.4 2.0% accuracy and 0.898 0.022 AUROC. fastprop takes only 1 minute and 54 seconds (1 minute and 39 seconds for descriptor calculation) and is competitive with the reference study in performance, achieving a 88.2 3.7% accuracy and 0.935 0.034 AUROC. The purely linear QSPR model falls far behind these methods with a 71.8 6.6% accuracy and 0.824 0.052 AUROC.
@@
First assembled and fitted to by Saldana and coauthors [35] the dataset (Flash) includes around 632 entries, primarily alkanes and some oxygen-containing compounds, and their literature-reported flash point. The reference study reports the performance on only one repetition, but manually confirms that the distribution of points in the three splits follows the parent dataset. The split itself was a 70/20/10 random split, which is repeated four times for this study.
@@
Assembled by Das and coauthors [36] from a collection of other smaller datasets, this dataset maps 442 molecular structures to a unified-scale Yield Sooting Index (YSI), a molecular property of interest to the combustion community. The reference study performs leave-one-out cross validation to fit a per-fragment contribution model, effectively a training size of >99%, without a holdout set. Though this is not standard practice and can lead to overly optimistic reported performance, the results will be carried forward regardless. The original study did not report overall performance metrics, so they have been re-calculated for this study using the predictions made by the reference model as provided on GitHub 444Predictions are available at this permalink to the CSV file on GitHub.. For comparison fastprop and Chemprop use a more typical 60/20/20 random split and 8 repetitions. Results are summarized in Table LABEL:ysi_results_table.
@@
The fastprop framework is not without its drawbacks. The two subsequent sections explore in greater detail two specific cases where fastprop loses out to existing methods, but some general notes about out-of-distribution predictions and overfitting are also included here. Like all machine learning methods, fastprop is not intended to make predictions outside of its training feature space. The use of molecular descriptors, which can become out-of-distribution, may exacerbate this problem but fastprop can optionally winsorize the descriptors to counteract this issue. Additionally, hyperparameter optimization of machine learning models in cheminformatics has been known to cause overfitting [39], especially on small datasets. Users should be cautious when optimizing fastprop models and rely on defaults when possible.
@@
DeepDelta reported an RMSE of 0.830 0.023 at predicting differences, whereas a typical Chemprop model trained to directly predict property values was only able to reach an RMSE of 0.965 0.09 when evaluated on its capacity to predict property differences. fastprop is able to outperform Chemprop, though not DeepDelta, achieving an RMSE of 0.930 0.029 when using the same splitting procedure above. It is evident that delta learning is still a powerful technique for regressing small datasets.
For completeness, the performance of Chemprop and fastprop when directly predicting the unbound fraction are also compared to the original study by Esaki and coauthors. Using both cross validation and and external test sets, they had an effective training/validation/testing split of 0.64/0.07/0.28 which will be repeated 4 times here for comparison. They used mordred descriptors in their model but as is convention they strictly applied linear modeling methods. All told, their model achieved an RMSE of 0.53 averaged across all testing data. In only 39 seconds, of which 31 are spent calculating descriptors, fastprop far exceeds the reference model with an RMSE of 0.207 0.024. This also surpasses Chemprop, itself outperforming the reference model with an RMSE of 0.223 0.036.
@@
Though not discussed here for the sake of length, fastprop contains the functionality to perform feature importance studies on trained models. By using SHAP values [28] once can assign a scalar ‘importance’ to each of the input features with respect to the target value, such as molecular weight having a significant positive impact on boiling point in alkanes. Experts users can leverage this information to guide molecular design and optimization or inform future lines of inquiry. Via the fastprop CLI users can train a model and then use fastprop shap to analyze the resulting trained network. fastprop shap will then generate diagrams to visualize the SHAP values.
@@ Acknowledgements
pre 4. Todeschini R, Consonni V (2009) Molecular descriptors for chemoinformatics. In: Methods and principles in medicinal chemistry. p 1252
@@
pre 5. Arockiaraj M, Paul D, Clement J, et al (2023) Novel molecular hybrid geometric-harmonic-zagreb degree based descriptors and their efficacy in QSPR studies of polycyclic aromatic hydrocarbons. SAR and QSAR in Environmental Research 34:569–589. https://doi.org/10.1080/1062936x.2023.2239149
@@
pre 6. Ma J, Sheridan RP, Liaw A, et al (2015) Deep neural nets as a method for quantitative structure-activity relationships. Journal of Chemical Information and Modeling 55:263–274. https://doi.org/10.1021/ci500747n
@@
pre 7. Coley CW, Barzilay R, Jaakkola TS, et al (2017) Prediction of organic reaction outcomes using machine learning. ACS Central Science 3:434–443. https://doi.org/10.1021/acscentsci.7b00064
@@
pre 8. Yang K, Swanson K, Jin W, et al (2019) Analyzing learned molecular representations for property prediction. Journal of Chemical Information and Modeling 59:3370–3388. https://doi.org/10.1021/acs.jcim.9b00237
@@
pre 9. Heid E, Greenman KP, Chung Y, et al (2024) Chemprop: A machine learning package for chemical property prediction. Journal of Chemical Information and Modeling 64:9–17. https://doi.org/10.1021/acs.jcim.3c01250
@@
pre 10. Song Y, Zheng S, Niu Z, et al (2021) Communicative representation learning on attributed molecular graphs. In: Proceedings of the twenty-ninth international joint conference on artificial intelligence. Yokohama, Yokohama, Japan
@@
pre 11. Zhou G, Gao Z, Ding Q, et al (2023) Uni-mol: A universal 3D molecular representation learning framework. In: The eleventh international conference on learning representations
@@
pre 12. Chen J, Schwaller P (2023) Molecular hypergraph neural networks. https://doi.org/10.48550/arXiv.2312.13136
@@
pre 13. Zhao B, Xu W, Guan J, Zhou S (2023) Molecular property prediction based on graph structure learning. arXiv preprint arXiv:231216855
@@
pre 14. Wang Z, Jiang T, Wang J, Xuan Q (2024) Multi-modal representation learning for molecular property prediction: Sequence, graph, geometry. https://doi.org/10.48550/ARXIV.2401.03369
@@
pre 15. Zhu Y, Chen D, Du Y, et al (2024) Molecular contrastive pretraining with collaborative featurizations. Journal of Chemical Information and Modeling 64:1112–1122. https://doi.org/10.1021/acs.jcim.3c01468
@@
pre 16. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
pre 17. Karpov P, Godin G, Tetko IV (2019) Transformer-CNN: Fast and reliable tool for QSAR. https://doi.org/10.48550/ARXIV.1911.06603
@@
pre 18. Tilborg D van, Brinkmann H, Criscuolo E, et al (2024) Deep learning for low-data drug discovery: Hurdles and opportunities. ChemRxiv. https://doi.org/10.26434/chemrxiv-2024-w0wvl
@@
pre 19. Moriwaki H, Tian Y-S, Kawashita N, Takagi T (2018) Mordred: A molecular descriptor calculator. Journal of Cheminformatics 10:4. https://doi.org/10.1186/s13321-018-0258-y
pre 20. Falcon W, The PyTorch Lightning team (2019) PyTorch Lightning
@@
pre 21. Weininger D (1988) SMILES, a chemical language and information system. 1. Introduction to methodology and encoding rules. Journal of Chemical Information and Computer Sciences 28:31–36. https://doi.org/10.1021/ci00057a005
pre 22. Sosnin S, Karlov D, Tetko IV, Fedorov MV (2019) Comparative study of multitask toxicity modeling on a broad chemical space. Journal of Chemical Information and Modeling 59:1062–1072. https://doi.org/10.1021/acs.jcim.8b00685
pre 23. McGibbon M, Shave S, Dong J, et al (2023) From intuition to AI: evolution of small molecule representations in drug discovery. Briefings in Bioinformatics 25:bbad422. https://doi.org/10.1093/bib/bbad422
@@
pre 24. Comesana AE, Huntington TT, Scown CD, et al (2022) A systematic method for selecting molecular descriptors as features when training models for predicting physiochemical properties. Fuel 321:123836. https://doi.org/https://doi.org/10.1016/j.fuel.2022.123836
@@
pre 25. Wu J, Wang S, Zhou L, et al (2020) Deep-learning architecture in QSPR modeling for the prediction of energy conversion efficiency of solar cells. Industrial and Engineering Chemistry Research 59:18991–19000. https://doi.org/10.1021/acs.iecr.0c03880
@@
pre 26. Esaki T, Ohashi R, Watanabe R, et al (2019) Computational model to predict the fraction of unbound drug in the brain. Journal of Chemical Information and Modeling 59:3251–3261. https://doi.org/10.1021/acs.jcim.9b00180